One cubic centimeter (1.0 cm3) of water has a mass of 1.0 ? 10?3 kg. (b) Assuming biological substances are 100% water, estimate the mass of a (spherical) cell with a diameter of 1.0 µm. (d) Assuming biological substances are 100% water, estimate the mass of a fly. Take a fly to be roughly a cylinder 4 mm long and 2 mm in diameter. Please show how you got the answer. And the answer has to be in kg.
Answer #1

a) Mass of a spherical cell of water of diameter d is

M_{cell}=Volume\times density=\frac{4\pi (d/2)^3}{3}\times\rho

this is equal to

M_{cell}=\frac{4\pi (10^{-6}/2)^3}{3}\times10^{-3}=5.236\times 10^{-16}\,\,kg

b) The mass of the fly is

M_{fly}=Volume\times density=\pi (d/2)^2 h\times\rho

where d is the diameter and h is the length

M_{fly}=\pi (0.002/2)^2 \times 0.004\times 10^{-3}=1.256\times 10^{-5}\,\,kg